### v2

# Theoretically Achieving Continuous Representation of Oriented Bounding Boxes

Abstract
Considerable efforts have been devoted to Oriented Object Detection (OOD). However, one lasting issue regarding the discontinuity in Oriented Bounding Box (OBB) representation remains unresolved, which is an inherent bottleneck for extant OOD methods. This paper endeavors to completely solve this issue in a theoretically guaranteed manner and puts an end to the ad-hoc efforts in this direction. Prior studies typically can only address one of the two cases of discontinuity: rotation and aspect ratio, and often inadvertently introduce decoding discontinuity, e.g. Decoding Incompleteness (DI) and Decoding Ambiguity (DA) as discussed in literature. Specifically, we propose a novel representation method called Continuous OBB (COBB), which can be readily integrated into existing detectors e.g. Faster-RCNN as a plugin. It can theoretically ensure continuity in bounding box regression which to our best knowledge, has not been achieved in literature for rectangle-based object representation. For fairness and transparency of experiments, we have developed a modularized benchmark based on the open-source deep learning framework Jittor’s detection toolbox JDet for OOD evaluation. On the popular DOTA dataset, by integrating Faster-RCNN as the same baseline model, our new method outperforms the peer method Gliding Vertex by 1.13% mAP50 (relative improvement 1.54%), and 2.46% mAP75 (relative improvement 5.91%), without any tricks.
Figure 1: Examples of Discontinuity in OBB Representations. (a) Acute-angle Representation limits the rotation angle of OBBs inside a range of $\frac{\pi}{2}$ ($[-\frac{\pi}{4},\frac{\pi}{4})$ in this example). The red $\text{OBB}_{1}$ and the blue $\text{OBB}_{2}$ are similar, but their representations are significantly different. (b) Long-edge Representation determines the rotation angle $\theta$ by the long side and the x-axis. A slight disturbance in the aspect ratio of square-like OBBs will cause a huge change in their representation, which causes Aspect Ratio Discontinuity. (c) CSL [37] divides the rotation angle into several classifications (6 classifications in this figure). OBB between two classifications cannot be accurately represented, which brings DI. (d) GWD [40] denotes OBBs by Gaussian distribution. As the squares with different rotation angles can correspond to the same Gaussian, the orientation of decoded squares will be ambiguous.
1 Introduction
Object detection constitutes a fundamental task within the realm of computer vision. In conventional object detection scenarios [52], the commonplace approach involves the localization of objects using Horizontal Bounding Boxes (HBB). However, in many real-world settings such as remote sensing [32, 3] and scene text [19, 13], where objects exhibit arbitrary orientations, HBBs are inadequate in precisely delineating object boundaries. To overcome this issue, Oriented Bounding Boxes (OBB) [32], conceptualized as rotated rectangles, have been introduced as a more suitable representation for Oriented Object Detection (OOD).
Various models have been proposed for OOD [2, 20, 17, 7, 35]. However, as illustrated in Figs. 1(a)-1(b), prevalent representation methods for OBBs exhibit discontinuity issues, encoding similar OBBs into distinct vectors. This introduces challenges in training neural networks as regression targets for similar input features may differ significantly, potentially causing confusion and hindering the training process. The relationship between two OBBs can be conceptualized as one being transformed into the other through geometric operations: translation, rotation, scaling, and aspect ratio changes. While translation and scaling are relatively benign, rotation and aspect ratio changes are primary sources of discontinuity.
Rotation discontinuity, often referred to as the “Boundary Problem” [37] or “Rotation Sensitive Error” [23], stems from the periodicity of rotation angles. Although prior efforts [37, 39, 46] have addressed this, most of them still suffer from discontinuity arising from changes in aspect ratio. Some other methods, e.g. Gliding Vertex [35], effectively address aspect ratio discontinuity. Nonetheless, these techniques continue to face challenges in overcoming rotation discontinuity. The discontinuity phenomena emerge during the encoding of OBBs into the regression target. Consequently, they can be termed encoding discontinuity.
Additionally, existing methods aim at resolving the encoding discontinuity issue, yet meanwhile would often bring about decoding discontinuity, namely Decoding Incompleteness (DI) and Decoding Ambiguity (DA). DI arises when OBBs cannot be accurately represented, often attributed to angle discretization and classification as exemplified by CSL [37] in Fig. 1(c). DA, on the other hand, pertains to instances where distinct OBBs share similar representations, rendering predicted OBBs sensitive to minor disturbances in the model’s output, exemplified by GWD [40] in Fig. 1(d). Fundamentally, DI and DA result in decoded OBBs lacking continuity concerning their representation. Hence, we categorize DI and DA as decoding discontinuities. The decoding discontinuities bring precision errors and directly degrade the prediction precision.
Due to the absence of a rigorous definition, prior approaches have often addressed discontinuity issues incompletely. To address this, we introduce formal continuity metrics, evaluating previous methods using these benchmarks. As a comprehensive solution to discontinuity problems, we propose Continuous OBB (COBB)—a novel, continuous OBB representation satisfying all defined metrics. COBB employs nine parameters derived from continuous functions based on the outer Horizontal Bounding Box (HBB) and OBB area. This ensures continuity as the outer HBB and OBB area undergo continuous changes during shape transformations. Our COBB can be easily integrated into existing OOD methods by simply replacing their original representations of OBB with ours.
We have developed a benchmark using the detection toolbox JDet of Jittor [10] which is an open-source deep learning framework friendly to vision tasks. In particular, a fair comparison across different models is made by aligning the data augmentation schemes and diverse techniques.
Our experiments on this benchmark demonstrate the effectiveness of COBB across diverse datasets and baseline detectors, particularly demonstrating advantages in high-precision object detection. Notably, it achieves a 3.95% improvement in $\text{mAP}_{75}$ when applied to Faster R-CNN on the DOTA Dataset. Detailed results are presented in Sec. 5.
Our contributions encompass the following aspects:
  • We systematically analyze the inherent discontinuity issues in existing OBB representation methods for OOD, and introduce formal metrics to assess their continuity.
  • Building upon our findings, we introduce COBB, a fully continuous representation of OBBs.
  • We construct a new benchmark for fair comparisons among OOD methods. Experiments on this benchmark validate the effectiveness of our approach, highlighting its advantages in high-precision OOD.
2 Related Work
2.1 Oriented Object Detection
With the increasing adoption of deep learning in computer vision, object detection models [16, 24, 30, 53, 5, 6, 11] have emerged to enhance computers’ capacity for recognizing objects in natural images. Typically tailored for predicting HBBs, these models serve as the foundation for OOD when augmented with modules for OBB prediction. Rotated Faster R-CNN [25] stands as a prominent baseline for OOD, replacing HBB regression targets with OBBs. Several OOD methods, such as RoI Transformer [2], Gliding Vertex [35], ReDet [7], and Oriented R-CNN [33], follow a similar structure and can be implemented on the Rotated Faster R-CNN framework.
While many OOD models share structural similarities, detailed implementation differences exist (e.g. Gliding Vertex [35] using ResNet101 as the backbone network, whereas CSL [37] employs ResNet50). To facilitate fair comparisons, we established a uniform pipeline with modular alternatives for implementing these models, minimizing implementation disparities.
2.2 Discontinuity in Oriented Object Detection
Methods aiming to handle the discontinuous representation of OBBs fall into three categories: Loss Improvement, Angle Encoding, and New OBB representation.
Loss Imporvement. Modifying the loss is a direct way to mitigate sudden changes in loss values caused by encoding discontinuity. Approaches like RIL [21] and RSDet [23] propose loss functions that approach zero as the model’s output converges to various representations of the ground truth OBB. PIoU [1] and SCRDet [38] incorporate Intersection over Union (IoU) between prediction results and regression targets in their loss. GWD [40], KLD [41], and KFIoU [44] convert OBBs into Gaussian distributions for IoU calculation, introducing potential DA for square-like objects. While showing empirical effectiveness in reducing the impact of discontinuity, these approaches do not provide a theoretical resolution to the problem.
Angle Encoding. Several methods focus on addressing the Periodicity of Angular (PoA), a primary cause of encoding discontinuity [37]. CSL [37] discretizes the rotation angle into a heavy regression target, with subsequent improvements by DCL [39], GF_CSL [29], MGAR [28], and AR-CSL [48]. While these methods enhance rotation continuity, most of them struggle with square-like objects and may introduce DI. PSC [46], FSTC [49], and ACM [34] encode the rotation angle into a continuous vector, yet they still exhibit discontinuity for square-like objects.
New OBB Representation. Other approaches explore alternative representations for OBBs instead of rectangles and rotation angles. Gliding Vertex [35] slides the four vertices of a HBB to construct an OBB. O2D-Net [31] and BBAVectors [45] represent an OBB using its center point and vectors from the center point to midpoints of its sides. PolarDet [51] and CRB [47] leverage polar coordinates, yet the rotation discontinuity still exists. DHRec [22] represents OBBs with double horizontal rectangles but struggles with distinguishing symmetrical tilted thin OBBs.
To the best of our knowledge, no method achieves perfect elimination of discontinuity. Previous approaches either fail in specific boundary situations or introduce DI and DA. The proposed COBB in this paper provides the first completely continuous representation of OBBs.
3 Theoretically Continuous Representation
In this section, we first introduce our devised metrics to assess the continuity of existing methods in Sec. 3.1. While Sec. 3.2 unveils our COBB that theoretically ensures continuity under these metrics. The continuity of COBB is rigorously demonstrated in Sec. 3.4, with comprehensive details provided in the supplemental material.
Table 1: Comparison of Methods Dedicated to Discontinuity. Tar (R), Tar (A), Loss (R), and Loss (A) stand for Target Rotation Continuity, Target Aspect Ratio Continuity, Loss Rotation Continuity, and Loss Aspect Ratio Continuity, respectively. Dec (C) and Dec (R) stand for Decoding Completeness and Decoding Robustness. Acute-Angle is the common OBB representation that limits the rotation angle into a range of $\frac{\pi}{2}$ ($[-\frac{\pi}{4},\frac{\pi}{4})$ for example). Long-Edge is another common OBB representation, which determines the rotation angle $\theta$ by the long side and the $x$-axis, and the $\theta$ is within a range of length $\pi$. Further explanation of this table is provided in the supplemental material, see Sec. 8.
Method Tar (R) Tar (A) Loss (R) Loss (A) Dec (C) Dec (R)
Acute-Angle - -
Long-Edge - - - -
RIL [21] - -
RSDet ($l_{mr}^{5p}$[23] - -
PIoU [1] - -
SCRDet [38] - -
GWD [40] -
KLD [41] -
KFIoU [44] -
CSL [37] - - - -
DCL [39] - - -
GF_CSL [29] - -
MGAR [28] - - -
AR-CSL [48] - - - -
PSC [46] - -
FSTC [49] - -
ACM [34] - -
Gliding Vertex [35] - - -
O2D-Net [31] - -
BBAVectors [45] - - -
PolarDet [51] - -
DHRec [22] -
CRB [47] - - -
Ours
3.1 Metrics for Continuity
As depicted in Fig. 1, prevalent methods for OBB prediction commonly face the challenge of encoding discontinuity. Prior endeavors often address some specific boundary cases, such as nearly horizontal OBBs and square-like OBBs [35, 46, 37]. While these methods may exhibit continuity in certain boundary scenarios, they often overlook others. For instance, CSL [37] maintains rotation continuity for nearly horizontal OBBs but fails for square-like OBBs. Furthermore, some methods assert the resolution of the discontinuity, yet they still struggle with sudden changes in regression targets [46, 34].
To formally define continuity, we introduce $f_{Enc}$ as the mapping function from an OBB to $R^{n}$, and $f_{Dec}$ as the reverse mapping from a subset of $R^{n}$ to an OBB. Notably, $R(x,\theta)$ denotes the transformation generating an OBB $y$ by rotating the initial OBB $x$ by $\theta$ in a clockwise direction. Meanwhile, $A(x,r)$ generates a set of OBBs $\{y,z\}$ by adjusting one side of OBB $x$ to be $r$ times its original length. We refer to $S$ as the set of OBBs, and $L$ symbolizes the loss function. Models utilize $f_{Enc}$ to convert OBBs into regression targets and employ $f_{Dec}$ to translate prediction results into estimated OBBs.
Target Rotation Continuity: Minor rotations should minimally affect the regression target.
$\small\forall x\in S,\lim_{\theta\to 0}||f_{Enc}(x)-f_{Enc}\left(R(x,\theta)\right)||=0.$ (1)
The corresponding discontinuity is often referred to as the “Boundary Problem” [37, 39].
Target Aspect Ratio Continuity: Slight changes in aspect ratio should minimally impact the regression target.
$\small\forall x\in S,\lim_{r\to 1}\sum_{y\in A(x,r)}||f_{Enc}(x)-f_{Enc}(y)||=0.$ (2)
Loss Rotation Continuity: Small rotations should minimally affect the loss value.
$\small\forall x\in S,\lim_{\theta\to 0}||L\left(f_{Enc}(x),f_{Enc}\left(R(x,\theta)\right)\right)||=0.$ (3)
Loss Aspect Ratio Continuity: Minor aspect ratio changes should minimally alter the loss value.
$\small\forall x\in S,\lim_{r\to 1}\sum_{y\in A(x,r)}||L\left(f_{Enc}(x),f_{Enc}(y)\right)||=0.$ (4)
Decoding Completeness: Every OBB can be accurately represented.
$\small\forall x\in S,\exists t\in R^{n},IoU\left(x,f_{Dec}(t)\right)=1.$ (5)
Decoding Completeness is equivalent to avoiding Decoding Incompleteness (DI) illustrated in Fig. 1(c).
Decoding Robustness: Decoded OBBs should be robust to slight errors in their representation.
$\forall x\in S,\forall\epsilon>0,\exists\xi>0,\forall\Delta d\in R^{n}\land||\Delta d||<\xi,$ (6)
$1-IoU\left(x,f_{Dec}\left(f_{Enc}(x)+\Delta d\right)\right)<\epsilon.$
Decoding Robustness is equivalent to avoiding Decoding Ambiguity (DA) illustrated in Fig. 1(d).
Previous research has covered the first four metrics, which we formally defined, whereas there’s been limited exploration of the last two metrics. Our investigation into existing OBB representation methods helped unveil the neglected discontinuity, which formed the basis for these two metrics. A further detailed explanation is provided in the supplemental material. Tab. 1 summarizes existing methods addressing discontinuity. However, these methods are not universally continuous. To comprehensively resolve the problem of discontinuity, we propose COBB, which ensures both encoding continuity and decoding continuity.
Figure 2: Example of COBB. COBB utilizes the outer HBB ($x_{c},y_{c},w,h$), sliding ratio $r_{s}$, and four IoU scores. (a) Example of the outer HBB and $r_{s}$. In this instance, $r_{s}=\frac{y_{2}-y_{1}}{h}$ when $w>h$, where $y_{1}$ and $y_{2}$ denote the two smaller y-coordinates among the four vertices of the OBB. (b) Using $x_{c}$, $y_{c}$, $w$, $h$, and $r_{s}$, along with the properties of similar triangles, we can derive and solve a system of equations to obtain the parameters for four OBBs (details provided in the supplemental material). Distinguishing between these OBBs is guided by the positional relationship between their vertices and the midpoints on each side.
3.2 Our Continuous Representation for OBB
Note that the outer HBB and the area of an OBB undergo continuous changes during shape transformations. Consequently, we sought to represent an OBB with a 5-dimensional vector, $(x_{c},y_{c},w,h,r_{a})$. Here, $(x_{c},y_{c})$, $w$, and $h$ refer to the center point, width, and height of the outer HBB, respectively, while $r_{a}$ is the acreage ratio of the OBB relative to its outer HBB.
It can be proven that only a pair of symmetrical OBBs shares the same $(x_{c},y_{c},w,h,r_{a})$ (detailed proof is provided in the supplemental material). However, directly computing OBBs from $(x_{c},y_{c},w,h,r_{a})$ is a complex process. To address this challenge, we introduce a sliding ratio, $r_{s}$, to estimate $r_{a}$, defined as follows.
$\small r_{s}=\begin{cases}\frac{x_{2}-x_{1}}{w}&w<h,\\ \frac{y_{2}-y_{1}}{h}&w\geq h,\\ \end{cases}$ (7)
where the x-coordinates of four vertices of the OBB are sorted as $x_{1},x_{2},x_{3},x_{4}$ from small to large, and y-coordinates are sorted as $y_{1},y_{2},y_{3},y_{4}$. It can be proved that $r_{s}$ can be computed as $r_{s}=f(\min(r_{a},1-r_{a}))$, where $f:[0,0.5]\rightarrow[0,0.5]$ is a continuous strictly increasing map (proof provided in the supplemental material). This implies the $r_{s}$ changes continuously as OBBs transform.
However, as shown in Fig. 2, a total of four different OBBs can be encoded into the same $(x_{c},y_{c},w,h,r_{s})$, leading to potential DA. To mitigate DA, we utilize Intersection over Unions (IoUs) between the target OBB and the four OBBs as scores for classification. Importantly, these IoUs can be directly computed using $(x_{c},y_{c},w,h,r_{s})$ and the classification of the target OBB, eliminating the need for complex computations involving IoU between arbitrary OBBs. The detailed computation process is provided in the supplemental material.
Finally, $(x_{c},y_{c},w,h,r_{s},s_{0},s_{1},s_{2},s_{3})$ will be considered as a continuous representation of OBBs, where $s_{0}$, $s_{1}$, $s_{2}$, $s_{3}$ are IoUs between the target OBB and the four OBBs with the same $(x_{c},y_{c},w,h,r_{s})$.
By reversing the above process, a 9-dimensional vector is decoded into a single OBB. Without loss of generality, assuming $w\geq h$, exploiting the properties of similar triangles allows the computation of $x_{2}-x_{1}$ and $y_{2}-y_{1}$:
$y_{2}-y_{1}$ $=r_{s}h,$ (8)
$x_{2}-x_{1}$ $=\frac{1-\sqrt{1-4\cdot\frac{h^{2}}{w^{2}}\cdot r_{s}(1-r_{s})}}{2}w.$
The classification with the highest IoU score determines the style of the generated OBB. Using the HBB, $x_{2}-x_{1}$, $y_{2}-y_{1}$, and the style, the coordinates of the OBB’s four vertices can be easily computed. The detailed computation process is provided in supplemental material.
3.3 Implementing COBB in OOD Models
Most models use the bias between the ground truth and the assigned proposal as the regression target to take advantage of the priori information of proposals. In our method, for horizontal proposal region $(x_{p},y_{p},w_{p},h_{p})$, the regression target is computed as follows, same as Faster R-CNN [25]:
$t_{x}$ $=\frac{\Delta x}{w_{p}},\ t_{y}=\frac{\Delta y}{h_{p}},\ t_{w}=\ln\left(\frac{w}{w_{p}}\right),\ t_{h}=\ln\left(\frac{h}{h_{p}}\right),$ (9)
where $box_{t}=(t_{x},t_{y},t_{w},t_{h})$ is the target for the outer HBB, and $\Delta x=x_{c}-x_{p}$, $\Delta y=y_{c}-y_{p}$.
According to Eq. 7, the value of $r_{s}$ lies within the range [0, 0.5]. To take advantage of this property, an effective way is to limit the range of prediction results, such as employing the sigmoid function. In this situation, the regression target for $r_{s}$ is computed as follows:
$\small r_{sig}=2r_{s}.$ (10)
Another method is extending the domain of $r_{s}$ as follows:
$\small r_{ln}=\begin{cases}1+\log_{2}(r_{s})&r_{a}<0.5,\\ 1+\log_{2}(1-r_{s})&r_{a}\geq 0.5.\end{cases}$ (11)
Compared with $r_{sig}$, $r_{ln}$ exhibits increased sensitivity to $r_{s}$ when $r_{a}$ is small, aiding detectors in precisely predicting inclined thin objects. Based on the definition of the regression target of $r_{s}$, our methods fall into two classifications: COBB-sig and COBB-ln.
The regression target of IoU scores is defined as:
$\small s_{t}=(s_{0}^{\lambda},s_{1}^{\lambda},s_{2}^{\lambda},s_{3}^{\lambda}),$ (12)
where $\lambda$ is a predefined constant to amplify the gap between scores of the ground truth and other classifications.
For models employing oriented proposal regions, we rotate the proposal region and target OBB around the center of the proposal region until its rotation angle becomes zero. Subsequently, we calculate the regression target as that for horizontal proposal regions. By reversing this process, OBBs can be easily recovered from the regression target and the oriented proposal regions. The detailed computation method is provided in the supplemental material.
In our approach, the loss function is defined as:
$L=$ $w_{1}L_{cls}+w_{2}L_{box}(box_{p},box_{t})$ (13)
$+w_{3}L_{r}(r_{p},r_{t})+w_{4}L_{s}(s_{p},s_{t}),$
where $box_{p}$, $r_{p}$, and $s_{p}$ denote predicted outer HBBs, predicted $r_{s}$, and predicted IoU scores, respectively. $r_{t}$ is the regression target of $r_{s}$, which is either $r_{sig}$ for COBB-sig or $r_{ln}$ for COBB-ln. $L_{cls}$ stands for the classification loss, which aligns with that of the baseline model (e.g. cross-entropy loss for Faster R-CNN [25]). $L_{box}$, $L_{r}$, and $L_{s}$ are Smooth L1 Loss [25]. The hyperparameters $w_{1}$, $w_{2}$, $w_{3}$, and $w_{4}$ are predefined constants.
3.4 Theoretical Guarantee of the Continuity
The COBB, as detailed in Sec. 3.2, theoretically ensures continuity under the metrics outlined in Sec. 3.1. Here, we briefly elucidate the reasons behind the continuity, with detailed proofs provided in the supplemental material.
Theoretical Analysis on Encoding Continuity: According to Sec. 3.2, $x_{c}$, $y_{c}$, $w$, $h$, and $r_{s}$ exhibit continuity concerning the outer HBB and the area of the target OBB. For unambiguous classifications, IoU scores remain continuous concerning $x_{c}$, $y_{c}$, $w$, $h$, and $r_{s}$. In cases of ambiguous OBB classifications, the IoU scores remain similar regardless of the classification. Consequently, the regression target produced by our method maintains continuity for both rotation and aspect ratio changes.
Theoretical Analysis on Decoding Continuity: The OBB generation process in Sec. 3.2 ensures precise reversal from the 9 parameters, mitigating inherent DI errors.
To avoid DA, the decoder must resist slight changes in its input. When IoU scores are fixed, the decoded four vertices remain continuous concerning $(x_{c},y_{c},w,h,r_{s})$. Notably, when $w$ is similar to $h$, the changes in the ordering of values between $w$ and $h$ do not lead to DA, as $x_{2}-x_{1}$ is similar to $y_{2}-y_{1}$. If $x_{c}$, $y_{c}$, $w$, $h$, and $r_{s}$ are fixed, and a slight perturbation in IoU scores results in a classification error, the IoU between the OBB before and after perturbation decoding is close to 1, adhering to the definition of IoU scores. In summary, our method exhibits resistance to perturbations in predicted results, thereby avoiding DA.
3.5 Further Comparison with Peer Methods
Compared with Gliding Vertex: The Gliding Vertex method [35] represents an OBB by sliding the four vertices of its outer HBB. However, rotation continuity is compromised when the OBB is nearly horizontal. Moreover, its decoded results manifest as irregular quadrilaterals, and refining these into accurate OBBs introduces accuracy errors. In contrast, our methods ensure continuous prediction targets and loss values for nearly horizontal OBBs, and the decoded quadrilaterals consistently represent accurate OBBs.
Compared with CSL-based methods: CSL-based methods [37, 39, 29, 28, 48] discretize rotation angles, converting angle regression into an angle classification problem to address rotation discontinuity. However, angle discretization introduces DI problems and results in a heavy prediction layer. Additionally, most CSL-based methods do not maintain continuity in aspect ratio changes when dealing with square-like OBBs. In contrast, our method ensures encoding continuity in both rotation and aspect ratio changes without introducing DI. Furthermore, our approach encodes an OBB using only 9 parameters.
Table 2: Open source OOD benchmarks.
Benchmark ArialDet OBBDet AlphaRotate [42] MMRotate [54] JDet
DL library PyTorch PyTorch Tensorflow PyTorch Jittor
Algorithm 5 10 18 19 20
Dataset 1 6 11 4 6
4 Benchmarking OOD under JDet
4.1 Brief Description of JDet
Our benchmark utilizes the Jittor object DETection models library (JDet), an open-source library dedicated to object detection, particularly supporting OOD methods. Built on Jittor [10], a deep learning framework, JDet facilitates the entire training and evaluation processes of object detection models. Preprocessing of diverse datasets precedes training or testing, ensuring a unified format. Various data augmentations, such as rotation and category balancing, are implemented as interchangeable and combinable modules. During testing, JDet supports diverse post-processing techniques for different datasets, with VOC2012 [4] serving as the implementation for evaluation. The library accommodates common object detection frameworks (e.g., Faster R-CNN [25]) and operators for OOD (e.g., RRoI Align [2]).
In total, JDet comprises 20 models and supports 6 datasets. A comparison between JDet and other open-source libraries is presented in Tab. 2.
4.2 Components for Unified Benchmarking
To mitigate variations between models, we categorized several modules and constructed OOD models by assembling these modules. The identified modules include:
  • Backbone: Extracts features from input images; most models employ ResNet [9] as the backbone network with FPN [14] for feature extraction at different scales.
  • Anchor Generation: Defines anchors for every pixel in the feature map.
  • Ground Truth Assignment: Assigns ground truth bounding boxes to proposal regions based on their IoU.
  • Result Generation Network: Neural networks for classifying anchors or proposal regions and regressing targets from regions.
  • Encoder/Decoder: Converts proposal regions into regression targets and model outputs into detection results.
  • Region of Interest Feature Extraction: Extracts features of proposal regions for further detection and refinement, focusing on object-level details compared to image-level backbone features.
  • Loss Function: Most models employ cross-entropy loss for classification and L1 loss for OBB prediction.
Implementing these modules consistently enhances the uniformity and comparability of the benchmarked models.
Table 3: mAP of models in JDet benchmark on DOTA-v1.0.
Model Venue $\text{mAP}_{50}$ $\text{mAP}_{75}$ $\text{mAP}_{50:95}$
H2RBox [43] ICLR’23 67.62 35.48 36.67
CSL [37] ECCV’20 67.99 34.51 36.43
RSDet [23] AAAI’21 68.41 36.93 37.91
RetinaNet [15] ICCV’17 68.18 36.84 38.15
KLD [41] NeurIPS’21 68.75 38.68 39.29
KFIoU [44] ICLR’23 68.99 35.00 37.59
GWD [40] ICML’21 69.02 38.48 39.62
FCOS [27] ICCV’19 70.37 39.78 40.25
ATSS [50] CVPR’20 72.44 39.81 41.08
[…]
Faster R-CNN [25] NeurIPS’15 73.01 40.13 41.33
Gliding Vertex [35] TPAMI’20 73.31 41.62 41.57
RoI Trans. [2] CVPR’19 75.59 48.54 46.35
Oriented R-CNN [33] ICCV’21 75.11 47.48 45.20
ReDet [7] CVPR’21 76.38 50.83 47.08
Ours (RoI Trans. based) - 76.53 50.41 46.97
Ours (ReDet based) - 76.52 51.38 47.67
4.3 Detection Models in the Benchmark
Experiments are conducted on multiple models within our benchmark framework, all subject to uniform conditions. The baseline models chosen for comparison were Rotated Faster R-CNN [25] and Rotated RetinaNet [15]. To minimize implementation discrepancies, most of the other models are implemented with minimal alterations to their corresponding baseline architectures.
For fairness, we standardized data processing and training settings across different experiments, following the detailed settings outlined in Sec. 5.1. The experimental results on DOTA-v1.0 are presented in Tab. 3, with additional results available in the supplemental material.
5 Experiments
5.1 Datasets and Implementation Details
DOTA [32] is a dataset for remote sensing object detection. We evaluated models on DOTA-v1.0 and DOTA-v1.5. DOTA-v1.0 comprises 2,806 aerial images whose resolution is between 800$\times$ 800 and 4,000$\times$ 4,000, and a total of 188,282 target instances are annotated, covering 15 common categories. DOTA-v1.5 maintains the same image and dataset segmentation as DOTA-v1.0 but introduces labeling for extremely small objects (less than 10 pixels) and incorporates the container crane (CC) category.
Figure 3: Visual results of KLD [41] and ours. Due to DA, KLD struggles to accurately predict the orientation of square-like objects. In contrast, our COBB circumvents DA, enhancing its precision in predicting the orientation of square-like objects.
Table 4: Comparison: IoU Scores vs. One-hot Coding. The experimental setup remains consistent with Rotated Faster R-CNN + COBB-sig, except for the OBB classification scores. All experiments were performed on DOTA-v1.0.
Scores $\text{mAP}_{50}$ $\text{mAP}_{75}$ $\text{mAP}_{50:95}$
One-hot 73.46 43.76 42.90
IoU scores 74.00 44.03 43.29
Table 5: Comparison of COBB across Proposal Types. Experiments were performed on DOTA-v1.0 using RoI Transformer, a model that incorporates both Horizontal Proposals (HPs) and Oriented Proposals (OPs), as the baseline.
HPs OPs $\text{mAP}_{50}$ $\text{mAP}_{75}$ $\text{mAP}_{50:95}$
- - 75.59 48.54 46.35
COBB-ln - 76.27 50.23 47.06
- COBB-ln 76.10 48.11 46.32
COBB-ln COBB-ln 76.53 50.41 46.97
Table 6: Comparing Different Regression Targets. The experimental setup mirrors that of Rotated Faster R-CNN + COBB-ln. One approach utilizes $r_{a}$, representing the OBB’s acreage ratio concerning its outer HBB, for Regression Target (RT) calculation, while the alternative method uses $r_{s}$, the sliding ratio. All experiments were conducted on DOTA-v1.0.
RT $\text{mAP}_{50}$ $\text{mAP}_{75}$ $\text{mAP}_{50:95}$
Using $r_{a}$ 74.13 43.31 42.94
Using $r_{s}$ 74.44 44.08 43.53
Table 7: mAP across datasets. COBB-sig takes $r_{sig}$ for horizontal proposals, and COBB-ln takes $r_{ln}$ for horizontal ones. COBB-ln-sig takes $r_{ln}$ for horizontal ones, and $r_{sig}$ for rotated ones. The definition of COBB-ln-ln and COBB-sig-sig is similar.
Models DOTA-v1.0 DOTA-v1.5 DIOR HRSC2016 FAIR1M-1.0 FAIR1M-2.0
$\text{mAP}_{50}$ $\text{mAP}_{75}$ $\text{mAP}_{50:95}$ $\text{mAP}_{50}$ $\text{mAP}_{75}$ $\text{mAP}_{50:95}$ $\text{mAP}_{50}$ $\text{mAP}_{75}$ $\text{mAP}_{50}$ $\text{mAP}_{75}$ $\text{mAP}_{50}$ $\text{mAP}_{50}$
Rotated Faster R-CNN [25] 73.01 40.13 41.33 63.52 35.36 35.96 60.64 35.26 83.34 31.64 35.16 40.16
Gliding Vertex [35] 73.31 41.62 41.57 63.12 36.98 36.32 61.49 36.24 92.23 58.52 36.37 40.82
+COBB-sig 74.00 44.03 43.29 64.03 36.88 37.17 62.28 37.70 92.69 68.87 36.81 41.11
+COBB-ln 74.44 44.08 43.53 64.35 37.62 37.30 62.58 37.55 92.71 72.29 36.53 41.23
RoI Trans. [2] 75.59 48.54 46.35 65.69 41.76 40.36 66.09 44.26 96.73 88.76 39.31 43.93
+COBB-sig-sig 76.49 50.26 46.63 65.88 42.76 40.85 66.72 45.01 96.72 90.60 39.61 44.42
+COBB-ln-sig 76.55 49.91 46.68 67.18 41.75 40.80 67.47 45.51 96.71 90.89 39.82 44.78
+COBB-ln-ln 76.53 50.41 46.97 66.66 43.29 40.96 67.53 45.27 97.19 91.35 39.66 44.54
Oriented R-CNN [33] 75.11 47.48 45.20 65.47 40.35 39.31 64.38 41.19 96.61 86.49 38.30 42.90
+COBB-sig 75.52 48.35 45.61 66.25 41.34 40.04 65.65 42.78 96.77 87.43 38.81 43.31
+COBB-ln 76.25 48.48 45.92 66.18 41.42 40.01 65.42 42.19 96.74 88.23 38.83 43.43
DIOR [3] serves as a large-scale resource for remote sensing object detection, encompassing a total of 23,463 images, and spanning 20 distinct target categories. As stipulated in [3], DIOR is partitioned into a training set of 11,725 images and a testing set of 11,738 images.
HRSC2016 [18] is a ship detection dataset. Our training incorporates both the training and validation sets, while the test set is reserved for assessing model accuracy.
FAIR1M [26], designed for fine object detection in aerial images, consists of 5 categories with 37 subcategories. The dataset is available in two versions, -1.0 and -2.0. For FAIR1M-1.0, model training utilized the training set, and model evaluation was performed on the test set. For FAIR1M-2.0, models were trained on both the training and validation sets, with evaluation conducted on the test set.
All experiments were performed using a single NVIDIA RTX 3090. The models utilized ResNet-50 [9] and FPN [14] to extract multi-level feature maps. SGD optimization was employed during the training stage. Data augmentation included random flipping, with each image having a 50% chance of horizontal flipping followed by a 50% chance of vertical flipping.
5.2 Ablation Study
Comparison between IoU Scores and One-hot Coding: In Sec. 3.2, we implemented IoU scores to differentiate OBBs sharing the same $(x_{c},y_{c},w,h,r_{s})$. Alternatively, one-hot coding seems simpler for classification. We compared the two methods on rotated Faster R-CNN + COBB-sig, as recorded in Tab. 4. The model’s accuracy using one-hot coding is lower than that using IoU scores due to the discontinuity introduced by one-hot coding.
Comparison of COBB across Proposal Types: The implementation of COBB on horizontal and oriented proposals is discussed in Sec. 3.3. To validate its effectiveness on both types, we conducted experiments on RoI Transformer [2], which employs both horizontal and oriented proposals. Results in Tab. 5 demonstrate integrating COBB enhances $\text{mAP}_{50}$ for both proposal types, with a more significant improvement observed in horizontal proposals.
Comparison between $r_{a}$ and $r_{s}$: In Sec. 3.2, we approximated $r_{a}$ with $r_{s}$. Further emphasizing the superiority of $r_{s}$ over $r_{a}$, experiments were conducted on Faster R-CNN, as shown in Tab. 6. The results illustrate that $r_{s}$ is more effective than $r_{a}$. This effectiveness stems from the complexity of recovering an OBB from the outer HBB and $r_{a}$, potentially leading to precision loss. Moreover, slight prediction errors on $r_{a}$ may cause significantly larger errors in the predicted OBB than errors caused by slight $r_{s}$ errors. Detailed insights are available in the supplemental material.
5.3 Results and Analysis
Detailed results on different datasets and detectors are presented in Tab. 7, with comprehensive ablation study details available in the supplemental material.
Results on DOTA: The results on DOTA-v1.0 show that Gliding Vertex outperforms Faster R-CNN by 0.30% in $\text{mAP}_{50}$ and 1.49% in $\text{mAP}_{75}$. Despite its accuracy advantage, Gliding Vertex suffers from discontinuity and DI, limiting its overall accuracy.
Our methods demonstrate superiority over existing approaches, especially in high-precision detection. Specifically, COBB outperforms Faster R-CNN by 1.21%, 3.92%, and 2.08% in $\text{mAP}_{50}$, $\text{mAP}_{75}$, and $\text{mAP}_{50:95}$ on average. On RoI Transformer and Oriented R-CNN, our method also significantly enhances the accuracy. For RoI Transformer, it improves $\text{mAP}_{50}$, $\text{mAP}_{75}$, and $\text{mAP}_{50:95}$ by 0.93%, 1.69%, and 0.41%, respectively. On Oriented R-CNN, it outperforms by 0.77%, 0.94%, and 0.56%, respectively. Notably, our method exhibits clear advantages in $\text{mAP}_{75}$, signifying its capability in high-precision object detection, a result of continuous representation and avoidance of DI and DA.
On DOTA-v1.5, our method effectively boosts the baseline detector accuracy. On average, the performance gain is 1.89%, 0.84%, and 1.03% in $\text{mAP}_{75}$ for Faster R-CNN, RoI Transformer, and Oriented R-CNN, respectively. This highlights its effectiveness in small object detection.
Results on HRSC2016: HRSC2016 involves ship objects that are in large aspect ratios. As shown in Tab. 7, a substantial gap is observed between COBB-sig and COBB-ln. As discussed in Sec. 3.3, COBB-ln’s advantage lies in its better capture of slight changes in $r_{s}$ for inclined large aspect ratio objects. Accordingly, COBB-ln outperforms COBB-sig by 3.42% in $\text{mAP}_{75}$ and Gliding Vertex by 13.77%. On RoI Transformer, COBB-ln-ln outperforms RoI Transformer by 2.59%.
Results on DIOR and FAIR1M: Experiments on less common datasets namely DIOR and FAIR1M, are also conducted. On DIOR, COBB-ln outperforms Faster R-CNN and RoI Transformer by 1.94% and 1.44%, respectively, in $\text{mAP}_{50}$, and by 2.29% and 1.01% in $\text{mAP}_{75}$. On FAIR1M-1.0 and FAIR1M-2.0, our method significantly improves baseline detectors.
Visualization Results: Fig. 3 visually compares the results of KLD [41] and COBB. KLD’s precision for the orientation of square-like objects is compromised by DA, whereas COBB accurately represents these objects, achieving strong performance by eliminating DA.
6 Conclusion
We have extensively shown the presence of boundary discontinuity in existing OOD models. To solve this problem, we have introduced COBB, an innovative continuous OBB representation method. Our experimental results showcase the effectiveness of our proposed method, achieving a notable improvement of 3.95% in $\text{mAP}_{75}$ on Rotated Faster R-CNN applied to the DOTA Dataset, without employing any additional techniques. COBB also has limitations. The outer HBB, sliding ratio $r_{s}$, and IoU scores exhibit irregular variations during OBB rotation, restricting its impact on rotation-equivariant detectors (e.g. ReDet [7]). Despite this, COBB proves effective in enhancing most OOD models by eliminating discontinuity.
7 Ackownledgement
The work was supported by the National Science and Technology Major Project under Grant 2021ZD0112902, the National Natural Science Foundation of China under Grant (62220106003, 62222607, 72342023), and the Research Grant of Beijing Higher Institution Engineering Research Center and Tsinghua-Tencent Joint Laboratory for Internet Innovation Technology.
Supplementary Material
8 Futher Explain of Metrics for Continuity
Within Sec. 3.1, we introduced diverse metrics for evaluating methodological continuity. These metrics encompass Target Rotation Continuity, Target Aspect Ratio Continuity, Loss Rotation Continuity, Loss Aspect Ratio Continuity, Decoding Completeness, and Decoding Robustness. While formal definitions were provided in Sec. 3.1, this section delves deeper into their conceptual underpinnings.
Target Rotation Continuity: This metric assesses whether each OBB is encoded into a sole prediction target, and if slight rotations induce gradual changes in the prediction target. Notably, PSC [46] demonstrates target rotation continuity by utilizing phase-shifting coding, ensuring continuous encoding despite OBB orientation changes. Conversely, Gliding Vertex [35] exhibits notable deviations in target rotation continuity when minor rotations affect nearly horizontal OBBs, leading to abrupt changes in the prediction target.
Target Aspect ratio Continuity: Here, the focus lies on determining if every OBB is encoded into a single prediction target and whether slight aspect ratio adjustments cause sudden changes in the prediction target. For instance, Gliding Vertex [35] maintains target aspect ratio continuity. However, PSC [46] struggles to sustain aspect ratio continuity, particularly when dealing with square-shaped OBBs.
Loss Continuity: Loss Continuity encompasses two distinct components: Loss Rotation Continuity and Loss Aspect Ratio Continuity. This metric evaluates whether minor rotation or aspect ratio changes result in abrupt fluctuations in the loss value. While some methods might falter in maintaining target rotation or aspect ratio continuity, they compensate by refining the loss function to ensure loss continuity. Notably, employing an L1 Loss function aids in preserving loss continuity for methods demonstrating target continuity.
Decoding Completeness: This criterion mandates precise representation of all OBBs. Methods rooted in CSL paradigms, such as those discussed in [37, 39, 29], often fall short in achieving Decoding Completeness due to discretized angle predictions, leading to imprecise orientation estimations of OBBs within finite angle classifications. Notably, we consider methods based on Gaussian distribution (such as GWD [40], KLD [41], and KFIoU [44]) satisfying Decoding Completeness because squares in various orientations can be possibly precisely decoded theoretically. However, in actual implementation, squares in only one orientation can be precisely decoded.
Decoding Robustness: Decoding Robustness demands that decoded OBBs remain resilient to slight errors in their representation. An example illustrating dissatisfaction with decoding robustness is GWD [40], which assigns square-like OBBs in different orientations to the same Gaussian distribution, leading to imprecise predictions for square-like targets. Especially, DHRec [22] encodes two symmetrical tilted slender OBBs into comparable representations, particularly when the aspect ratio is significantly large. Although this is uncommon, it suggests a propensity for the algorithm to confuse slender OBBs that are oriented in differing directions. Consequently, this observation leads us to conclude that DHRec does not fulfill the criteria for Decoding Robustness.
Target Rotation Continuity, Target Aspect ratio Continuity, Loss Rotation Continuity, and Loss Aspect Ratio Continuity are collectively referred to as Encoding Continuity, while Decoding Completeness and Decoding Robustness are collectively referred to as Decoding Continuity.
9 Details of COBB
9.1 Derivation of Four OBBs with Identical Outer HBB and $r_{s}$
Figure 4: An OBB with Its outer HBB.
We establish the outer HBB ($x_{c},y_{c},w,h$) of an OBB along with the sliding ratio $r_{s}$ in Sec. 3.2. Herein, we elaborate on the existence of precisely four OBBs sharing the same outer HBB and $r_{s}$.
Given a generic OBB characterized by the outer HBB $x_{c},y_{c},w,h$, and sliding ratio $r_{s}$, we employ the principles of similar triangles to derive the equation:
$||P_{1}B||\cdot||P_{2}B||=||P_{1}A||\cdot||P_{2}C||.$ (14)
Here, $||P_{1}B||$, $||P_{2}B||$,$||P_{1}A||$, and $||P_{2}C||$ are line segments depicted In Fig. 4. When $w\geq h$, $||P_{1}B||\cdot||P_{2}B||=h^{2}r_{s}(1-r_{s})$. Conversely, when $w<h$, $||P_{1}A||\cdot||P_{2}C||=w^{2}r_{s}(1-r_{s})$. Assuming $w\geq h$, we deduce:
$\begin{cases}||P_{1}B||\cdot(h-||P_{1}B||)=h^{2}r_{s}(1-r_{s}),\\ ||P_{1}A||\cdot(w-||P_{1}A||)=h^{2}r_{s}(1-r_{s}).\\ \end{cases}$ (15)
Solving these equations yields:
$\begin{cases}||P_{1}B||=\frac{1\pm(1-2r_{s})}{2}h,\\ ||P_{1}A||=\frac{1\pm\sqrt{1-4\cdot\frac{h^{2}}{w^{2}}\cdot r_{s}(1-r_{s})}}{2}w.\end{cases}$ (16)
These solutions delineate four distinct groups, corresponding to four unique OBBs. The detailed process for constructing these OBBs is elucidated in Sec. 9.2.
9.2 OBB Recovery from Nine Parameters
This section elucidates the method for computing the OBB from its outer HBB, sliding ratio $r_{s}$, and IoU scores $s_{0}$, $s_{1}$, $s_{2}$, and $s_{3}$.
Building upon the methodology outlined in Sec. 9.1 for computing two OBB vertices using Eq. 16, the four OBBs can be derived as follows:
$\begin{split}OBB_{0}=&[(x_{c}-x_{s},y_{c}-0.5h),(x_{c}+0.5w,y_{c}+y_{s}),\\ &(x_{c}+x_{s},y_{c}+0.5h),(x_{c}-0.5w,y_{c}-y_{s})],\\ OBB_{1}=&[(x_{c}+x_{s},y_{c}-0.5h),(x_{c}+0.5w,y_{c}+y_{s}),\\ &(x_{c}-x_{s},y_{c}+0.5h),(x_{c}-0.5w,y_{c}-y_{s})],\\ OBB_{2}=&[(x_{c}-x_{s},y_{c}-0.5h),(x_{c}+0.5w,y_{c}-y_{s}),\\ &(x_{c}+x_{s},y_{c}+0.5h),(x_{c}-0.5w,y_{c}+y_{s})],\\ OBB_{3}=&[(x_{c}+x_{s},y_{c}-0.5h),(x_{c}+0.5w,y_{c}-y_{s}),\\ &(x_{c}-x_{s},y_{c}+0.5h),(x_{c}-0.5w,y_{c}+y_{s})].\\ \end{split}$ (17)
Here, $y_{s}$ and $x_{s}$ are computed as:
$y_{s}$ =$=\begin{cases}\frac{1-2r_{s}}{2}h&w\geq h,\\ \frac{\sqrt{1-4\frac{w^{2}}{h^{2}}r_{s}(1-r_{s})}}{2}h&w<h,\end{cases}$ (18)
$x_{s}$ =$=\begin{cases}\frac{\sqrt{1-4\frac{h^{2}}{w^{2}}r_{s}(1-r_{s})}}{2}w&w\geq h,\\ \frac{1-2r_{s}}{2}w&w<h.\end{cases}$
The OBB associated with the highest IoU score is selected as the recovered result.
9.3 COBB Implementation for Oriented Proposals
Figure 5: COBB for Oriented Proposals.
Many models utilize the discrepancy between ground truth and assigned proposals as an n-dimensional regression target. In Sec. 3.3, we introduced employing COBB to compute the bias between the ground truth OBB and a horizontal proposal region. Here, we extend this method to oriented proposal regions.
As depicted in Fig. 5, for an oriented proposal with center $(x_{c},y_{c})$ and rotation angle $\theta$ (clockwise), the regression target computation involves the following steps:
  • Rotate the oriented proposal and ground truth OBB by $\theta$ counterclockwise around the center $(x_{c},y_{c})$.
  • Calculate the bias between the rotated proposal region and ground truth, leveraging the method introduced in Sec. 3.3 for horizontally oriented regions.
For predicting the OBB from an oriented proposal and predicted vector, the process involves:
  • Rotate the oriented proposal and ground truth OBB by $\theta$ counterclockwise around the center $(x_{c},y_{c})$.
  • Compute based on the rotated horizontally oriented proposal and the predicted bias to obtain an OBB.
  • Rotate the OBB calculated in the previous step by $\theta$ clockwise to derive the predicted OBB.
This methodology suits Oriented Object Detection (OOD) models using Rotated RoI Align for feature extraction from oriented regions [2, 33]. However, if a model utilizes a different method for feature extraction from oriented regions, the proposed regression target generation method may not be applicable.
9.4 Relationship between $r_{s}$ and $r_{a}$
We assert that the sliding ratio $r_{s}$ provides an approximation for $r_{a}$, the acreage ratio of the OBB concerning its outer HBB in Sec. 3.2. In this section, we establish the relationship between $r_{s}$ and $r_{a}$ and subsequently demonstrate why only a pair of symmetrical OBBs share identical $(x_{c},y_{c},w,h,r_{a})$.
In Sec. 9.1, we established that only four OBBs share the same $(x_{c},y_{c},w,h,r_{s})$. For any OBB, we express $r_{s}$ in terms of $r_{a}$ and derive the unified equation:
$4r_{s}(1-r_{s})=\frac{r_{wh}^{2}+1-\sqrt{(r_{wh}^{2}+1)^{2}-16r_{wh}^{2}r_{a}(1-r_{a})}}{2r_{wh}^{2}}.$ (19)
Here, $r_{wh}=\min(\frac{w}{h},\frac{h}{w})$ denotes the aspect ratio of the outer HBB. Eq. 19 indicates that $r_{s}(1-r_{s})$ monotonically increases with respect to $r_{a}(1-r_{a})$. Since four OBBs share identical $x_{c},y_{c},w,h,r_{s}$, they also share the same $r_{a}(1-r_{a})$. Given that $r_{a}$ for $OBB_{0}$ and $OBB_{3}$ is below $0.5$, while for the other two, it exceeds $0.5$, we conclude that only a symmetrical OBB pair shares identical $(x_{c},y_{c},w,h,r_{a})$.
9.5 Details of Computing IoU scores
Directly computing IoU scores often involves generating four OBBs based on $(x_{c},y_{c},w,h,r_{s})$ and then evaluating the IoU between these OBBs and the ground truth. However, this direct calculation can be intricate and time-consuming. Our approach involves first computing the IoUs between the four OBBs sharing identical $(x_{c},y_{c},w,h,r_{s})$ based on the five parameters.
The simplest method for computing IoU scores is directly generating the four OBBs according to $(x_{c},y_{c},w,h,r_{s})$, and then calculating the IoU between the ground truth and the four OBBs. However, the calculation of IoU between OBBs is complex and time-consuming. We notice that the IoU between the four OBBs sharing the same $(x_{c},y_{c},w,h,r_{s})$ can be directly computed from the five parameters. As a result, we can compute these IoUs first, and then select the IoU scores based on the type of the ground truth OBB.
Assuming $w\geq h$, let’s define intermediate variables:
$r_{sx}$ $=\frac{1-\sqrt{1-4\cdot\frac{h^{2}}{w^{2}}\cdot r_{s}(1-r_{s})}}{2},r_{sy}=r_{s},$ (20)
$l_{1}$ $=\sqrt{(r_{sx}w)^{2}+(r_{sy}h)^{2}},$
$l_{2}$ $=\sqrt{[(1-r_{sx})w]^{2}+[(1-r_{sy})h]^{2}},$
$l_{3}$ $=\sqrt{(r_{sx}w)^{2}+[(1-r_{sy})h]^{2}},$
$l_{4}$ $=\sqrt{[(1-r_{sx})w]^{2}+(r_{sy}h)^{2}},$
where $l_{1}$, $l_{2}$, $l_{3}$, and $l_{4}$ are correspond to potential side lengths of the OBBs.
IoU between $OBB_{0}$ and $OBB_{1}$, $IoU_{0,1}$ is:
$\begin{split}I_{0,1}&=[1-\frac{(1-2r_{sx})r_{sx}w^{2}}{(1-r_{sy})h^{2}}]l_{1}l_{2},\\ IoU_{0,1}&=\frac{I_{0,1}}{l_{1}l_{2}+l_{3}l_{4}-I_{0,1}}.\end{split}$ (21)
IoU between $OBB_{0}$ and $OBB_{2}$, $IoU_{0,2}$ is:
$\begin{split}I_{0,2}&=(1-\frac{(1-2r_{sy})r_{sy}h^{2}}{(1-r_{sx})w^{2}})l_{1}l_{2},\\ IoU_{0,2}&=\frac{I_{0,2}}{l_{1}l_{2}+l_{3}l_{4}-I_{0,2}}.\end{split}$ (22)
IoU between $OBB_{0}$ and $OBB_{3}$, $IoU_{0,3}$ is:
$\begin{split}I_{0,3}&=\frac{(r_{sx}+r_{sy}-2r_{sx}r_{sy})^{2}}{(1-r_{sx})(1-r_{sy})}\times\frac{wh}{2},\\ IoU_{0,3}&=\begin{cases}\frac{I_{0,3}}{2l_{1}l_{2}-I_{0,3}}&I_{0,3}\neq 0,\\ 0&I_{0,3}=0.\end{cases}\end{split}$ (23)
IoU between $OBB_{1}$ and $OBB_{2}$, $IoU_{1,2}$ is:
$\begin{split}h_{1}&=\frac{1}{2}w-\frac{\frac{1}{2}-r_{sy}}{1-r_{sy}}r_{sx}w,\\ h_{2}&=\frac{1}{2}h-\frac{\frac{1}{2}-r_{sx}}{1-r_{sx}}r_{sy}h,\\ \tan\alpha&=\frac{\frac{\frac{1}{2}-r_{sx}}{1-r_{sx}}l_{4}}{\frac{1}{2(1-r_{sy})}l_{3}},\quad\tan\beta=\frac{\frac{\frac{1}{2}-r_{sy}}{1-r_{sy}}l_{3}}{\frac{1}{2(1-r_{sx})}l_{4}},\\ I_{1,2}&=2\frac{\tan\alpha\tan\beta}{\tan\alpha+\tan\beta}(h_{1}^{2}+h_{2}^{2})+2h_{1}h_{2},\\ IoU_{1,2}&=\begin{cases}\frac{I_{1,2}}{2l_{3}l_{4}-I_{1,2}}&\tan\alpha\tan\beta\neq 0,\\ \frac{2h_{1}h_{2}}{2l_{3}l_{4}-2h_{1}h_{2}}&\tan\alpha\tan\beta=0.\end{cases}\end{split}$ (24)
These IoUs constitute the IoU matrix $M(w,h,r_{s})$ as:
$M(w,h,r_{s})=\begin{bmatrix}1&IoU_{0,1}&IoU_{0,2}&IoU_{0,3}\\ IoU_{0,1}&1&IoU_{1,2}&IoU_{0,2}\\ IoU_{0,2}&IoU_{1,2}&1&IoU_{0,1}\\ IoU_{0,3}&IoU_{0,2}&IoU_{0,1}&1\\ \end{bmatrix}.$ (25)
This matrix ensures continuity for $w>0,h>0,r_{s}\in[0,0.5]$. Each element in row $i$ and column $j$ of $M(w,h,r_{s})$ represents the IoU between $OBB_{i}$ and $OBB_{j}$.
Given a ground truth OBB, we identify its corresponding type among these four OBBs and extract the corresponding row from $M(w,h,r_{s})$ as its IoU scores.
9.6 Proof of Encoding Continuity of COBB
In this section, we demonstrate the continuous evolution of the nine parameters in COBB as the OBB transforms.
The outer HBB and the area of an OBB change continuously during shape transformations, ensuring the continuity of $x_{c}$, $y_{c}$, $w$, $h$, and $r_{a}$. Eq. 19 substantiates the continuity of $r_{s}$ concerning $r_{a}$, thereby ensuring the continuity of $r_{s}$ as well.
To establish the continuity of IoU scores, we consider an OBB perturbed from OBB $X$, denoted as $Y$. The similarity between $x_{c}$, $y_{c}$, $w$, $h$, and $r_{s}$ of $X$ and $Y$ emerges from their continuous evolution during shape transformations. Eq. 17 confirms the analogous construction of four OBBs using the parameters of $X$ and $Y$. IoU scores represent the overlaps between $X$ and the four OBBs sharing the same $x_{c}$, $y_{c}$, $w$, $h$, and $r_{s}$ as $X$. As both $X$ and the associated four OBBs remain analogous before and after perturbation, minor disturbances on $X$ do not significantly alter its IoU scores.
Figure 6: COBB Parameters during OBB Transformation. AR represents the Aspect Ratio.
Figure 7: Traditional Acute-angle Representation Parameters during OBB Rotation. The traditional Acute-angle Representation encodes an OBB using the center $(x_{c},y_{c})$, width $w$, height $h$, and rotation angle $\theta$, which is constrained within the range $[-\frac{\pi}{4},\frac{\pi}{4})$.
Fig. 6 and Fig. 7 provide comparative insights into the regression targets of COBB and the traditional Acute-angle Representation. COBB consistently exhibits continuous encoding results, while the Acute-angle Representation displays distinct shifts at rotation angles of $\pi/4$ and $43\pi/$ 4.
In summary, slight disturbances to an OBB minimally affect its $x_{c}$, $y_{c}$, $w$, $h$, $r_{s}$, and IoU scores, affirming the stability of these parameters under minor perturbations.
9.7 Proof of Decoding Continuity of COBB
Decoding Completeness is inherently fulfilled within the COBB decoding process. Here, we aim to establish the Decoding Robustness of COBB.
Slight disturbances in the parameters of COBB can be categorized as perturbations in the outer HBB, $r_{s}$, and IoU scores. As per Eq. 17, minor perturbations in the outer HBB and $r_{s}$ won’t significantly alter the decoded OBB if IoU scores remain constant. If a perturbation doesn’t affect the classification corresponding to the highest IoU score, it won’t impact the decoded OBB. Even if a perturbation shifts the highest score from classification $i$ to $j$, where $s_{i}=1$ (indicating the correct classification), slight disturbances maintain the relative values of $s_{i}$ and $s_{j}$, making $s_{j}$ close to 1. As $s_{j}$ represents the IoU between the correct OBB and the decoded $OBB_{j}$ (from Eq. 17), choosing $OBB_{j}$ as the decoded result won’t introduce significant decoding errors.
9.8 Comparison of Different Regression Targets
To mitigate the sensitivity of objects with large aspect ratios to predicted results, we introduce $r_{ln}$ as the regression target, which is defined by Eq. 11. It is an approximation to $f_{ln}(r_{a})=1+\log_{2}(r_{a})$. The direct use of $f_{ln}(r_{a})$ is feasible by recovering $r_{s}$ with Eq. 19. Notably, only two OBBs share the same $x_{c}$, $y_{c}$, $w$, $h$, and $r_{ln}$; hence, we need to derive $r_{s}$ from $r_{ln}$ first and then select one of the four OBBs that share the same $r_{s}$ as the decoded result, preventing Decoding Ambiguity.
In previous experiments in Tab. 6, we established the superiority of $r_{ln}$ over $f_{ln}(r_{a})$. One reason is that the process of recovering $r_{s}$ from $r_{a}$ can introduce precision errors. This section presents another reason: the sensitivity of the decoded OBB to slight disturbances in $f_{ln}(r_{a})$.
Figure 8: Impact of Slight Disturbance on $r_{s}$ and $r_{a}$: We assessed the effects on the decoded OBB when $r_{s}$ and $r_{a}$ undergo slight disturbances. To mitigate the sensitivity of OBBs with large aspect ratios to small perturbations in $r_{s}$ or $r_{a}$, we consider $f(r_{s})=r_{ln}$ (as previously defined) and $f_{ln}(r_{a})=1+\log_{2}r_{a}$ as the specific perturbed values.
Table 8: NAE of parameters. $\alpha$ is the parameter for orientation determination in Gliding Vertex. In our method and Gliding Vertex xy denotes the center, wh represents width and height, $r_{s}$ and IoU scores are parameters of our method, and $\alpha$ denotes the additional parameters for orientation determination of Gliding Vertex.
Method xy wh $r_{s}$ IoU scores $\alpha$
Ours. 0.18% 0.31% 0.28% 0.41% -
Glidng Vertex 0.17% 0.31% - - 0.75%
Table 9: $\text{mAP}_{50}$ of Models in JDet Benchmark on DOTA-v1.0.
Model PL BD BR GTF SV LV SH TC BC ST SBF RA HA SP HC $\text{mAP}_{50}$
CSL 89.36 77.74 37.66 66.73 77.85 59.27 77.75 90.86 79.74 82.23 55.20 63.07 51.55 66.03 44.93 67.99
RsDet 89.47 78.02 38.39 63.01 77.28 60.95 76.41 90.77 81.80 83.47 54.26 61.98 51.68 66.55 52.18 68.41
RetinaNet 89.56 79.70 38.12 67.58 76.21 59.47 76.80 90.70 83.56 81.07 53.88 64.73 53.36 65.99 41.97 68.18
RetinaNet+KLD 89.42 76.51 39.36 65.20 77.83 63.19 82.48 90.90 79.20 83.46 54.39 63.83 53.26 67.67 44.53 68.75
RetinaNet+KFIoU 89.38 81.05 39.12 68.39 77.37 62.58 77.86 90.87 82.43 82.32 56.00 65.60 53.50 67.08 41.36 68.99
RetinaNet+GWD 89.10 78.04 39.07 69.21 77.27 62.05 81.05 90.90 84.31 83.02 57.24 61.96 53.79 64.40 43.83 69.02
FCOS 88.07 78.80 44.85 65.58 74.88 68.17 77.84 90.90 79.99 83.60 57.42 65.30 62.96 68.59 48.55 70.37
ATSS 88.44 78.77 49.14 67.17 77.63 74.97 85.56 90.78 83.39 83.82 59.17 61.39 65.35 65.48 55.50 72.44
[…]
Faster R-CNN 89.46 83.89 49.64 69.59 77.57 73.23 86.52 90.90 79.33 85.74 58.84 60.49 65.78 68.64 55.55 73.01
Gliding. 89.34 83.68 50.15 69.97 78.20 72.51 87.17 90.90 79.94 85.46 57.07 62.57 66.93 66.12 59.61 73.31
RoI Trans. 89.21 83.88 53.01 72.97 77.86 78.08 88.01 90.86 86.94 85.84 63.50 61.53 75.77 70.33 56.02 75.59
O-RCNN. 89.72 84.41 52.94 71.80 78.71 77.51 88.15 90.90 85.90 84.90 61.58 63.93 74.23 69.94 51.99 75.11
ReDet 88.90 82.28 52.42 72.76 77.63 82.52 88.12 90.88 86.51 85.81 67.34 65.33 76.07 68.58 60.54 76.38
Ours (RoI Trans. based) 89.52 84.98 54.99 72.16 77.71 82.81 88.10 90.81 85.45 85.62 63.89 66.15 76.64 70.13 59.05 76.53
Ours (ReDet based) 89.71 84.82 53.27 71.40 77.02 83.80 88.07 90.85 87.11 86.20 66.44 63.72 76.15 67.99 61.26 76.52
Table 10: mAP of COBB on DOTA-v1.0 under Multi-scale Data Augmentation.
Models $\text{mAP}_{50}$ $\text{mAP}_{75}$ $\text{mAP}_{50:95}$
Oriented R-CNN [33] 78.73 55.07 50.57
+COBB-sig 79.09 55.61 50.80
+COBB-ln 79.23 56.15 50.55
Table 11: Experiments on Optimizing IoU Enhancement Factor $\lambda$. We utilize $IoU^{\lambda}$ as a specific regression target for IoU scores to widen the distinction between scores for the ground truth and other classifications. These experiments were conducted using Rotated Faster R-CNN + COBB-sig.
$\lambda$ $\text{mAP}_{50}$ $\text{mAP}_{75}$ $\text{mAP}_{50:95}$
1 73.41 44.21 43.10
2 74.00 44.03 43.29
4 73.63 43.85 42.83
Table 12: Experiments on SOTA baselines on DOTA-v1.0.
Method $\text{mAP}_{50}$ $\text{mAP}_{75}$ $\text{mAP}_{50:95}$
SES 75.72 48.86 46.19
SES + Ours. 76.43(+0.71) 49.28(+0.42) 46.59(+0.40)
LSKNet-t 76.68 49.28 46.15
LSKNet-t + Ours. 77.29(+0.61) 50.91(+1.63) 47.62(+1.47)
Fig. 8 illustrates the impact of slight disturbances on $r_{ln}$ and $f_{ln}(r_{a})$. The impact is quantified as follows:
$\Delta OBB=(1-IoU[f_{Dec}(r),f_{Dec}(r+\epsilon)])\epsilon^{-1}.$ (26)
Here, $r$ is $r_{ln}$ or $f_{ln}(r_{a})$, $f_{Dec}$ decodes $r$ with a nearly square outer HBB into an OBB, and $\epsilon$ is a small value close to 0. Notably, when $f_{ln}(r_{a})$ approaches zero, the decoded OBB demonstrates significant sensitivity to variations in $f_{ln}(r_{a})$, whereas the impact of slight disturbances on $r_{ln}$ remains relatively stable.
9.9 Models’ Ability to Well Estimate Parameters
We contend that the parameters in our COBB are easily estimable due to their continuity. To elucidate this assertion, we introduce the Normalized Average Error (NAE) as a metric for assessing the difficulty of parameter estimation. Given the $i$-th prediction of the parameter as $P_{i}$ and its corresponding ground truth as $T_{i}$, the NAE is defined as $\textrm{NAE}=\frac{1}{N}\sum_{i=1}^{N}\frac{(P_{i}-T_{i})^{2}}{(\max(T)-\min(T))^{2}}$, where $N$ represents the number of predictions, and $\max(T)$ and $\min(T)$ denote the maximum and minimum values of ground truth values.
We posit that parameters with small NAE values are more ready to estimate. The NAE values of our method and Gliding Vertex are documented in Tab. 8. Without significantly influencing the prediction difficulty of the other parameters, $r_{s}$ and IoU scores in our method can be better estimated than $\alpha$s in Gliding Vertex.
10 Additional Experiments
10.1 Training Settings
Experiments were conducted using Jittor [10] on a single RTX 3090 running on Linux. The models utilized ResNet-50 [9] with FPN [14] to extract multi-level feature maps. During training, an SGD optimizer was employed, with a learning rate of 0.005 for two-stage models and 0.01 for one-stage models.
For dataset-specific training:
  • DOTA, FAIR1M, and DIOR datasets were trained for 12 epochs, while HRSC2016 was trained for 36 epochs.
  • For images in the DOTA and FAIR1M datasets larger than $024\times 1,$ 024 × 1 , 024, they were split into multiple $024\times 1,$ 024 × 1 , 024 tiles with a 200-pixel overlap.
  • Data augmentation included random horizontal and vertical flips, each with a $50\%$ probability.
10.2 More Results on JDet Benchmark
Tab. 9 provides a detailed breakdown of the results presented in Tab. 3. These comprehensive experimental findings underscore the superior performance achieved by our proposed method within our benchmark evaluation.
Analysis of the results indicates a pronounced advantage of our method in detecting objects characterized by a long aspect ratio, exemplified by categories such as Bridge (BR), Large Vehicle (LV), and Harbor (HA). This advantage is attributable to the inherent continuity embedded within our methodology, which mitigates potential confusion and interference during the training process arising from sudden changes in regression targets as OBBs approach a horizontal orientation. Notably, the models implemented in our benchmark employ the Acute-angle Representation as a default approach. The conspicuous discontinuity inherent in the Acute-angle Representation becomes especially evident when objects exhibit a considerable aspect ratio. Consequently, the discernible advantage exhibited by our proposed method in these scenarios underscores the efficacy of its continuous nature.
10.3 Futher Ablation Study for COBB
COBB Performance under Multi-scale Data Augmentation. Tab. 10 showcases the performance of COBB on DOTA-v1.0 under Multi-scale data augmentation. Multi-scale augmentation involved resizing training images to 0.5, 1.0, and 1.5 times their original dimensions, and these variations were incorporated into training and testing. For large images split into tiles, the width of the overlapping area was adjusted to 500 pixels. The results affirm the efficacy of COBB under this data augmentation technique.
Optimizing the IoU Enhancement Factor $\lambda$. We alter the IoU scores by exponentiating them to the power of $\lambda$ to diminish the impact of incorrect categories. Tab. 11 displays experiments demonstrating that $\lambda=2$ achieves optimal performance. Notably, $\lambda=1$ exhibits advantages in $\text{mAP}_{75}$. Throughout this article, COBB-sig employs $\lambda=2$, while COBB-ln uses $\lambda=1$.
10.4 Experiments on Latest Techniques
To verify the advantage of our method on SOTA methods, we added our method to the latest SOTAs, including SES [36] and LSKNet [12]. Experiments in Tab. 12 demonstrate that our method achieves a relatively large $\text{mAP}_{50}$ improvement, 0.71% and 0.61%, over SES and LSKNet, respectively.